# Multi-site Integrated Optical Addressing of Trapped Ions

### v2

@@ Abstract
Ions stored in RF Paul traps Paul:1953 have provided an effective platform for quantum information due to their stability wu:2021 , long coherence times wang:2021 , and high fidelities clark:2021 . In addition to computing srinivas:2021 ; leung:2018 , ion traps have also been used for atomic clocks brewer:2019a ; Burt:2021Nat , quantum networks moehring:2007 ; Nichol:2022Nat , quantum sensors marciniak:2022 , and fundamental science Zhang:2018NC ; Pinkas:2023NP . The development of surface traps chiaverini:2005 ; blain:2021 furthered the potential for advancing these applications by enabling scaling to more ions. Using microfabrication capabilities originally developed by the semiconductor industry, surface ion traps were fabricated that could support many-ion quantum computers pino:2021 ; noel:2022 . These traps also provided a convenient platform for integrating other electronic and optical technologies moody:2022 , which are necessary for the individual addressing and readout of growing numbers of ions. These technologies include photonic waveguides, waveguide splitters Vasques:2022PRL , grating couplers mehta:2016 ; niffenegger:2020 ; ivory:2021 , detectors setzer:2021 ; Reens:2022PRL ; todaro:2021 , and modulators hogle:2023 , all of which combined can support quantum systems with many ions as well as applications which require low size, weight, and power (SWaP).
While many of these proof-of-principle experiments focused on the operation of a single ion site, simultaneous independent control of multiple ions at several trap sites had until now not been realized. Here, we demonstrate a room temperature ion trap and simultaneous optical addressing of three ${}^{171}\textrm{Yb}^{+}$ions in independent wells, using only light delivered through waveguides. The light addressing two of the three ions comes from a single split input, forming a multi-ion ensemble. Multi-mode interferometer (MMI) splitters are employed to equally divide the light from the input waveguide into two separate waveguides, which are routed to output couplers individually addressing the trapped ions. Other splitting ratios are possible by varying the interaction length in the MMI splitters, as are other fanout ratios (e.g. 1$\times$×N). This technique could be employed to deliver light to a much larger number of ions, ultimately limited by the power handling capacity of waveguides. A demonstration of 300 mW of 729 nm light in a silicon nitride waveguide mehta:2020 shows that many sites could be supported from a single launch. Large collections of ions are particularly interesting for atomic clocks, where it has been shown that separately interrogated ensembles of ions can achieve sensitivities that scale as $(\alpha N)^{-m/2}$, where $N$ is the number of ions in each of $m$ ensembles, and $\alpha$ is a protocol-dependent constant borregaard:2013 .
@@ I Experimental setup
II Multi-ion trapping and waveguide characterization
State detection is particularly challenging in surface traps using integrated waveguides because light emitted from output couplers can scatter into the imaging system and overwhelm the ion fluorescence. In a standard detection scheme Streed:2011PRL ; niffenegger:2020 , ions fluoresce at the same wavelength as the detection laser, rendering wavelength filters too broad to distinguish between the two. The light output from a grating is highly polarized and therefore polarization filtering is possible, however it was found to be insufficiently discriminating in this experiment.
@@ II Multi-ion trapping and waveguide characterization
With the side-imaging system, the objective lens collected and focused light onto either a CCD camera (Andor Luca) or a 32 channel, linear PMT array (Hamamatsu H11659-200) with a multi-pinhole filter for the three waveguide sites. The 100 $\mu$m  pinholes spatially filtered the ion fluorescence from background scattering to improve the signal-to-noise ratio (SNR) on the PMTs. Ultimately this arrangement achieved SNRs between 5 and 10 while supporting detection times on the order of milliseconds (4 ms for the measurement in Fig. 3). Fig. 2a shows fluorescence images of the ions illuminated by both free-space and waveguide beams. Each waveguide-illuminated ion, labeled (i) to (iii), was separated by a distance of 200 $\mu$m and placed in a local potential minimum created by the surrounding electrodes. To trap multiple ions simultaneously, we loaded each ion sequentially. The first ion was loaded at the loading hole position (marked with a red circle in Fig. 2a) using free-space beams (Doppler and repump) with a single-site potential solution (see Methods). This loaded ion was then shuttled towards the far-right position (iii) by applying the appropriate voltage waveforms to the interior control electrodes. This shuttling process was carried out without the use of cooling beams between sites, but once the ion reached the target position it was cooled by light from the output gratings. We subsequently loaded new ions by applying an additional potential minimum and repeating this process until three ions were stored at locations (i) to (iii).
@@
III Simultaneous multi-ion/multi-state addressing
@@
IV Cross-talk characterization
Figure 4: Optical cross-talk a, Measured cross-talk between the single ion ensemble (orange) and two ion ensemble (light and dark blue). Rabi oscillations with the D-state ($m_{F}=-1$) were measured while addressing the single ion ensemble (orange). Cross-talk was evident in the slow transition rate of the nearer ion (light blue) in the two ion ensemble (light and dark blue), and was more attenuated for the more distant ion. b, A similar measurement as a, but instead addressing the two ion ensemble (light blue and dark blue). Cross-talk was evident in the slow transition rate of the single ion ensemble (orange). c, Relative cross-talk intensities of transition light. Ratios were calculated from fits of Rabi frequencies and plotted for each trap site.
@@ IV Cross-talk characterization
V TOWARD A FULLY-INTEGRATED SYSTEM
@@ Acknowledgements.
We thank the members of Sandia’s Microsystems and Engineering Sciences Application (MESA) facility for their fabrication expertise and for helpful comments on the manuscript. This work was supported by the Defense Advanced Research Projects Activity (DARPA). Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under Contract No. DE-NA0003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
@@
J.K., W.J.S., D.L.S. and H.J.M. conceived the experiment. J.K. and W.J.S. took the measurement and analyzed data. M.G. designed the integrated optics while also performing waveguide tests and simulations with N.K. . J.V.D.W. developed multi-ion solution. R.L. setup the Doppler laser system. J.K., W.J.S., D.S. and H.J.M. wrote the original manuscript. All the authors reviewed and edited the manuscript. H.J.M. supervised the project.

### v3

@@ Abstract
Ions stored in RF Paul traps Paul:1953 have provided an effective platform for quantum information due to their stability wu:2021 , long coherence times wang:2021 , and high fidelities clark:2021 . In addition to computing srinivas:2021 ; leung:2018 , ion traps have also been used for atomic clocks brewer:2019a ; Burt:2021Nat , quantum networks moehring:2007 ; Nichol:2022Nat , quantum sensors marciniak:2022 , and fundamental science Zhang:2018NC ; Pinkas:2023NP . The development of surface traps chiaverini:2005 ; blain:2021 furthered the potential for advancing these applications by enabling scaling to more ions. Using microfabrication capabilities originally developed by the semiconductor industry, surface ion traps were fabricated that could support many-ion quantum computers pino:2021 ; noel:2022 . These traps also provided a convenient platform for integrating other electronic and optical technologies moody:2022 , which are necessary for the individual addressing and readout of growing numbers of ions. These technologies include photonic waveguides, waveguide splitters Vasques:2022PRL , grating couplers mehta:2016 ; niffenegger:2020 ; mehta:2020 ; ivory:2021 , detectors setzer:2021 ; Reens:2022PRL ; todaro:2021 , and modulators hogle:2023 , all of which combined can support quantum systems with many ions as well as applications which require low size, weight, and power (SWaP).
While many of these proof-of-principle experiments focused on the operation of a single ion site, simultaneous independent control of multiple ions at several trap sites had until now not been realized. Here, we demonstrate a room temperature ion trap and simultaneous optical addressing of three ${}^{171}\textrm{Yb}^{+}$ions in independent wells, using only light delivered through waveguides. The light addressing two of the three ions comes from a single split input, forming a multi-ion ensemble. Multi-mode interferometer (MMI) splitters are employed to equally divide the light from the input waveguide into two separate waveguides, which are routed to output couplers individually addressing the trapped ions. Other splitting ratios are possible by varying the interaction length in the MMI splitters, as are other fanout ratios (e.g. 1$\times$×N). This technique could be employed to deliver light to a much larger number of ions, ultimately limited by the power handling capacity of waveguides. Large collections of ions are particularly interesting for atomic clocks, where it has been shown that separately interrogated ensembles of ions can achieve sensitivities that scale as $(\alpha N)^{-m/2}$, where $N$ is the number of ions in each of $m$ ensembles, and $\alpha$ is a protocol-dependent constant borregaard:2013 .
@@ I.1 Experimental setup
I.2 Multi-ion trapping and waveguide characterization
State detection is particularly challenging in surface traps using integrated waveguides because light emitted from output couplers can scatter into the imaging system and overwhelm the ion fluorescence. In a standard detection scheme Streed:2011PRL ; niffenegger:2020 , ions fluoresce at the same wavelength as the detection laser, rendering wavelength filters too broad to distinguish between the two. The light output from a grating is highly polarized and therefore polarization filtering is feasible, but provided insufficient filtering of scattered light in this experiment.
@@ I.2 Multi-ion trapping and waveguide characterization
With the side-imaging system, the objective lens collected and focused light onto either a CCD camera (Andor Luca) or a 32 channel, linear PMT array (Hamamatsu H11659-200) with a multi-pinhole filter for the three waveguide sites. The 100 $\mu$m  pinholes spatially filtered the ion fluorescence from background scattering to improve the signal-to-noise ratio (SNR) on the PMTs. Ultimately this arrangement achieved SNRs between 5 and 10 while supporting detection times on the order of milliseconds (4 ms for the measurement in Fig. 3). Fig. 2a shows fluorescence images of the ions illuminated by both free-space and waveguide beams. Each waveguide-illuminated ion, labeled (i) to (iii), was separated by a distance of 200 $\mu$m and placed in a local potential minimum created by the surrounding electrodes. To trap multiple ions simultaneously, we loaded each ion sequentially. The first ion was loaded at the loading hole position (marked with a red circle in Fig. 2a) using free-space beams (Doppler and repump) with a single-site potential solution (see Methods). This loaded ion was then shuttled towards the far-right position (iii) by applying the appropriate voltage waveforms to the interior control electrodes. This shuttling process was carried out without the use of cooling beams between sites, but once the ion reached the target position it was cooled by light from the output gratings. We subsequently loaded new ions by applying an additional potential minimum and repeating this process until three ions were stored at locations (i) to (iii) (see Supplementary Fig. 2).
@@
I.3 Simultaneous multi-ion/multi-state addressing
@@
Figure 4: Optical cross-talk a, Measured cross-talk between the single ion ensemble (orange) and two ion ensemble (light and dark blue). Rabi oscillations with the D-state ($m_{F}=-1$) were measured while addressing the single ion ensemble (orange). Cross-talk was evident in the slow transition rate of the nearer ion (light blue) in the two ion ensemble (light and dark blue), and was more attenuated for the more distant ion. b, A similar measurement as a, but instead addressing the two ion ensemble (light blue and dark blue). Cross-talk was evident in the slow transition rate of the single ion ensemble (orange). Error bars show the standard error. c, Relative cross-talk intensities of transition light. Ratios were calculated from fits of Rabi frequencies and plotted for each trap site.
I.4 Cross-talk characterization
@@
I.5 Toward a fully-integrated system
@@ II Discussion
III METHODS
Trap design and fabrication
This trap was designed and fabricated at Sandia National Laboratories and shares features common to previous traps with integrated photonics ivory:2021 ; setzer:2021 , notably waveguides and MMI splitters for optical addressing as well as SPADs for state readout. The SPADs could not be operated when light was delivered via integrated waveguides, as the scatter from those components overwhelmed the ion signal on the SPADs. We implemented a grounded metal mesh above the SPAD to reduce unpredictable charging effects ivory:2021 and consequent shuttling problems. It serves a similar function as an optically transparent and electrically conductive coating (e.g. indium tin oxide). The mesh allowed the ion to be transported over the entire aperture of the SPAD while the SPAD was active, which was not possible in previous Sandia-fabricated SPAD traps setzer:2021 . The mesh was designed to be 1 $\mu$m thick with many square 3 $\mu$m $\times$ 3 $\mu$m openings to allow ion fluorescence to pass through. (see Supplementary Fig. 1). The mesh transmits roughly 55% of the ion fluorescence to the SPAD, and based on electrostatic simulations reduces the electric field at the ion by 16 dB compared to not having a screen.
Experimental system
Ion trapping.
To avoid coating the electrodes with ytterbium, the trap has several loading holes that penetrate the entire thickness of the trap and its substrate. Ytterbium metal in a stainless steel tube is heated with current to vaporize atoms in the direction of the loading holes. The ytterbium atoms are ionized using a two-photon process with laser beams at 393 nm and 399 nm, at which point ytterbium ions are electrically confined by the RF and DC fields created by the trap electrodes.
The trap potential well is positioned 50 $\mu$m above the chip’s surface. In the loading zone ions are illuminated and cooled using free-space Doppler (369 nm) and repump (935 nm) beams. Additional sidebands of 2.1 GHz and 14.7 GHz are applied to the Doppler beam using electro-optic modulators (EOMs) for optical pumping and Doppler cooling. During the optical pumping sequence, the ion is prepared in the |$\ket{2\textrm{S}_{1/2},F=0,m_{\textrm{F}}=0}$ state by applying only the extra 2.1 GHz frequency to the standard Doppler frequency. This preparation can be observed through a modified detection step that only detects atoms in the $F=1$ state.
The motional heating rate $\dot{\bar{n}}$ of the axial trap mode was measured for an ion at site (ii) illuminated with both free-space and waveguide beams. For the free-space case, we used sideband thermometry roos:2000 on the |$\ket{2\textrm{S}_{1/2},F=0,m_{\textrm{F}}=0}$ to |$\ket{2\textrm{D}_{3/2},F=2,m_{\textrm{F}}=-1}$ transition, and measured a heating rate of $\dot{\bar{n}}$ = $2.71\pm 1.63$ q/ms for an ion that was initially cooled to $\bar{n}\sim 2$. Using waveguides, we measured a heating rate using the Rabi flopping method described below. In this case a free-space beam was used only for detection to allow for top-imaging. This enabled more accurate measurements due to the higher fidelity detection and shorter detection times, while still using the integrated waveguides for all other light delivery. As presented in Supplementary Fig. 3, we measured $\dot{\bar{n}}=1.25\pm 0.14$ q/ms, comparable to the case without any integrated light delivery.
Optical losses in waveguides.
Each waveguide on the device has both an edge coupler and grating coupler input that are combined on chip, though only the grating couplers were used for this work. The 369 nm and 435 nm waveguides were separately tested and a total power reduction of $\approx$20 dB for each was measured. This includes the power reduction from the input combiner (3 dB) and splitter (3 dB) for separate sites, as well as a total loss of $\approx$14 dB. This total loss is the sum of the estimated insertion loss at the input grating (6-7 dB), waveguide propagation loss (1.35 dB/cm at 369 nm and 0.9 dB/cm at 435 nm over 5 mm), and the simulated output grating efficiency (6-7 dB). While higher power reductions were measured on some devices, the optically good ones were consistently in the 20 dB range.
In experiments with ions, a total loss of 49 dB for the 435 nm light was calculated based on the input power and Rabi rate of the ion, significantly higher than the $\approx$14 dB ideal total loss measured outside the vacuum chamber. A 369 nm waveguide on the actual device used in the experiments measured $\approx$14 dB total loss prior to experiments, but the 435 nm waveguide was not measured, and may have had significantly higher loss due to fabrication imperfections. In the experiment each laser beam is launched 75 mm from the grating outside the vacuum chamber and passes through a viewport at a $45^{\circ}$ angle, causing significant beam distortion that may contribute to the different total loss. Future systems will employ edge coupling to improve coupling efficiency, as demonstrated in other experiments mehta:2020 and described below.
Waveguide scalability.
In addition to the $1\times 2$ splitter used in this paper, other splitting ratios are possible by varying the interaction length in the MMI splitters, as are other fanout ratios (1×1$\times$1 ×N). Larger values of N reduce the number of required optical launches to the chip but increase the amount of power in the input waveguide. This is generally not an issue for visible wavelengths; a demonstration of 300 mW of 729 nm light in a silicon nitride waveguide mehta:2020 shows that many sites could be supported from a single launch. However, UV wavelengths tend to cause greater damage, so the development of low-loss waveguides that do not degrade from photodarkening with UV powers up to $\sim$100 mW is critical to realizing this approach to scaling.
To show that much better coupling is possible for UV wavelengths, we edge-coupled 369 nm light from a lensed fiber to a test chip and back into a separate lensed fiber. The test chip used the same waveguide material and structure as the trap. Assuming equal coupling losses on and off chip and subtracting for the measured waveguide absorption loss (0.6 dB), we measured a 4.1 dB coupling loss from the fiber to the chip. Over short time periods the total power loss remained stable for in-waveguide powers ranging from 7 mW to 24 mW, but the power degraded over longer (hour) time periods. Multiple causes for this degradation are possible, including surface contamination at the fiber or edge faces, absorption in the inverse taper used for coupling, or photodarkening in the waveguide itself. Future research will be conducted to understand this behavior.
Side-imaging system.
One of the main difficulties in integrated waveguide light delivery schemes is detection, since scattered photons from the grating output couplers are at the same wavelength as the ion fluorescence and cannot be filtered. To address this a side-imaging technique was developed. The fluorescence from an ion was collected with a 2-inch diameter objective lens from the side of the chamber, and guided to either a CCD camera or linear PMT array. The bottom half of photons emitted by the ion are obscured by the chip, which gives an effective NA=0.12, compared to the conventional top imaging system with NA=0.28.
A 1D PMT was used for simultaneous independent detection of multiple ions. The ion separation in the image plane was comparable to the cell spacing of the PMT array so that each ion was detected by separate consecutive PMT pixels. To exclude scattered light from the trap, a pinhole array with 100 $\mu$m diameter holes was placed in front of the PMT array.
Generating multi-ion trapping solutions for multiple locations.
The DC voltage solutions for most ion trap experiments are created by constraining the electric field to be zero along three axes and the derivative of the electric field to have a specific positive value along one axis, all at the single point where we wish to trap the ion. This defines a potential well with a specific motional frequency. Since this problem is underconstrained and there are many possible solutions, other requirements (e.g. voltage limits on the electrodes) and desirements (e.g. applying voltages to fewer electrodes) are applied. Small changes to compensate the ion for deviations in the local electric field were computed separately and applied additively.
While this system served well for many uses, the need to precisely position multiple ions with respect to fixed laser beams, as well as the need to shuttle newly loaded ions into position, required a more flexible solution. Initially a strategy of pre-computing all combinations of positions for all wells was attempted. Since the number of possible combinations of well positions grows exponentially with the number of ions, so did our voltage solutions. This quickly became infeasible. The resolution to this problem was to create these solutions in real-time. For the real-time multi-ion solution the constraints were specified then combined to form a linear system that could be solved. The desired positions of all ions are provided as an input, along with the compensation fields at those positions. The system looks up the appropriate constraint matrices, concatenates them, and solves the resulting system.
Theoretical approaches
Rabi flopping in the Lamb-Dicke regime
To quantify the average number of motional quanta in the system, we fit the experimentally obtained Rabi flop data to a dephasing model.
First, we check if our system is in Lamb-Dicke regime ($\sqrt{\bar{n}_{k}}\eta_{k}\ll 1$). Here the Lamb-Dicke parameter $\eta_{k}$ is defined as $\eta_{k}=\frac{2\pi}{\lambda}\cos{\theta}\sqrt{\hbar/(2m\omega_{k})}$ where $\lambda=435$ nm is the wavelength, $m$ is the mass of ${}^{171}\textrm{Yb}^{+}$, $\hbar$ is plank’s constant, $\theta$ is the beam angle, and $\omega_{k}$ is the secular motional frequency. With the given parameters, we confirmed that $\sqrt{\bar{n}_{k}}\eta_{k}\ll 1$ if $\bar{n}<50$, which is valid in our experiment.
Because the depths of ion traps are much larger than the thermal energy of the ion, we fit the Rabi carrier function in a Lamb-Dicke regime as $P(t)$, the total population of excited state, as follows Semenin:2022OLP :
$P(t)=\frac{a}{2}\left(1-\frac{f_{1}(t)}{f_{2}(t)}\right)$ (1)
where
$f_{1}(t)=Re[e^{i\Omega t}]\prod_{k=1}^{N}e^{-i\Omega{\eta^{2}_{k}}t/2}\left(1-\frac{\bar{n}e^{i\Omega{\eta^{2}_{k}}t}}{\bar{n}+1}\right)$ (2)
and
$f_{2}(t)=\prod_{k=1}^{N}\left((\bar{n}+1)-2\bar{n}\cos{(\Omega_{0}\eta^{2}_{k}t)}+\frac{\bar{n}^{2}}{\bar{n}+1}\right),$ (3)
for generalized $N$ normal vibrational modes, characterized by $N$ secular frequencies $\omega_{k}$. Here $a$ is the initial population of the ground state, $\Omega_{0}$ is the Rabi frequency of the ion at rest, and $\eta_{k}$ is the Lamb-Dicke parameter for the given ion in the $k$th mode.
Cross-talk - evanescent coupling calculation
[…]
Data availability
All relevant data are available from the corresponding authors upon reasonable request.
@@ Acknowledgements.
We thank Justin T. Schultz for his contribution in the early stages of side-imaging system setup and Tharon D. Morrison for the waveguide UV tolerance test. We thank the members of Sandia’s Microsystems and Engineering Sciences Application (MESA) facility for their fabrication expertise and for helpful comments on the manuscript. This work was supported by the Defense Advanced Research Projects Activity (DARPA). Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under Contract No. DE-NA0003525. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
@@
J.K., W.J.S., D.S. and H.J.M. conceived the experiment. J.K. and W.J.S. took the measurement and analyzed data. M.G. designed the integrated optics while also performing waveguide tests and simulations with N.K. J.V.D.W. developed multi-ion solution. R.L. setup the Doppler laser system. M.B. oversaw the trap fabrication. J.K., W.J.S., D.S. and H.J.M. wrote the original manuscript. All the authors reviewed and edited the manuscript. H.J.M. supervised the project.
Competing Interests.
J.K., W.J.S., and H.J.M. are inventors in patent applications related to the "side-imaging" technique used to collect ion fluorescence. The remaining authors declare no competing interests.
IV Supplementary figures
Figure 5: Lithography mask in the SPAD region A metal mesh is located on top of the SPAD (purple circle), consisting of 3 $\mu$m $\times$ 3 $\mu$m holes with 1 $\mu$m wide metal traces. The mesh and SPAD are surrounded by nearby electrodes, also shown in Fig. 1a.
Figure 6: Multi-ion loading sequence Three waveguide-illuminated independent ions are loading sequentially, starting from one loading location that used free-space beams. 3+1 ions are trapped simultaneously by generating potential wells and loading/shuttling ions via steps (1) to (5). The potential plot at (5) represents a real four-ion solution that is used in this work, providing secular frequencies of $2\pi\times 1.02$MHz.
Figure 7: Heating rate measurement A heating rate of $1.25\pm 0.14$ q/ms is measured using fully integrated beams for state preparation, cooling, and Rabi flopping, while a free-space beam is used for detection. inset: Measured Rabi flopping for different wait times.